OpenStudy (anonymous):

how do i find the derivative of y=(x^3) / x-1. Do i use the quotient rule?

OpenStudy (anonymous):

Yep. For u/v, the derivative is \[\frac{ v(u')-u(v')}{ v^2 }\] with u' = derivative of u.

OpenStudy (anonymous):

i dont know if i did this right but i got (3x^2-x^3) / x-1

OpenStudy (phi):

in your case u is x^3 and v is (x-1) u' is 3x^2 and v' is 1 v u' is (x-1) * 3x^2= 3x^3 - 3x^2 u v' = x^3*1 = x^3 v u' - u v' = 3x^3 - 3x^2 - x^3 that simplifies to 2x^3 - 3x^2 so the final answer is \[ \frac{x^2(2x-3)}{(x-1)^2} \]

OpenStudy (anonymous):

so i wouldnt use the equation g(x)*f'(x)-f(x)*g'(x) / (g(x))^2 or does both equations give me the same ans

OpenStudy (phi):

I used that formula, but I used u and v for the names rather than f and g

OpenStudy (anonymous):

o ok